40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Mon Sep 29 23:28:37 2014, Evan, DailyProgress, PMC, Re-inserted PMC, moved BB EOM noiseBudget.pdf
    Reply  Tue Sep 30 21:45:14 2014, Evan, DailyProgress, PMC, Re-inserted PMC, moved BB EOM 
    Reply  Mon Dec 15 20:43:49 2014, Aidan, Kate, Evan, DailyProgress, PMC, South Faraday isolator needs help 
       Reply  Sat Jun 20 22:59:58 2015, Evan, DailyProgress, optic, South Faraday isolator path reworked southfi.jpg
Message ID: 1528     Entry time: Mon Sep 29 23:28:37 2014     Reply to this: 1529   1534
Author: Evan 
Type: DailyProgress 
Category: PMC 
Subject: Re-inserted PMC, moved BB EOM 

I swept the south laser with a triangle wave and optimized the mode-matching as best I could using the periscope mirrors and the translation stages. I got to a visibility of 0.3, which stinks (the maximum is 0.7 with these birefringent coatings).

I took a beat spectrum (attached) and noticed that the noise around 0.1–1 kHz is improved. Indeed, by reducing the visibility south I find the beat gets worse.

I decided some more involved mode-matching (involving beam profiling and alm simulation) is needed.

Before setting up the beam profiler, I noticed that the beam entering the cavity does not appear Gaussian, as seen on an IR card.

The laser beam entering the first Faraday isolator appears to be 1–2 mm too low. It is clipping on the input aperture, and the transmitted beam looks like crap.

Neither the Faraday nor the laser itself have any alignment adjustment knobs. I therefore had to choose between two evils: shim up the laser mount (and thereby risk having to realign the entire optical path, as well as possibly reducing the heatsinking of the mount to the table) or reinsert the PMC and move the BB EOM.

I opted for the latter: I reinserted the PMC, removed the EOAM (+QWP+PBS), placed the resonant EOM where the EOAM used to be, and then placed the BB EOM where the resonant EOM used to be.

I will optimize the alignment through these components, check the polarization, and then take a new beat.

Attachment 1: noiseBudget.pdf  1.088 MB  | Hide | Hide all
noiseBudget.pdf
ELOG V3.1.3-